OpenStudy (anonymous):

10 gallon bucket of product will produce 100 pounds which will fill 1 cubic yard. How much material we it take to fill 600 square feet area 2 inches?

jigglypuff314 (jigglypuff314):

so 10gal = 1yd^3 = 9ft^3 you are given 600ft^2 * 2in 2in = 1/6 ft so 600ft^2 * 1/6 ft = 100ft^3

jigglypuff314 (jigglypuff314):

to get 100ft^3 to gal. multiply by (10/9)